# Elevating Software Trust: Unveiling and Quantifying the Risk Landscape

### v1

@@ VIII-C Comparison
Table VI: Comparison - Proposed SRiQT Framework Risk with/without Self-Generated Parameters.
@@
Table VII: Risk Scores by OWASP Risk Calculator.
@@
Table VI presents the comparison of the risk scores calculated using the proposed SRiQT framework by first employing the entire dataset $\mathcal{D}$ and then $\mathcal{D}_{x}$ is used which excludes the parameters (i.e., regarding the date of the software, experience in language, developer’s and publisher’s experience, number of downloads, update frequency, context, and penalty/reward) for which self-generated data points were utilized. While computing the risk scores for $\mathcal{D}_{x}$, equation 11 was modified to remove the parameter for a number of downloads and normalize the value of user ratings to bring it into the range [0,1]. The values computed with all the parameters differ significantly from the ones calculated by employing only a subset of all the parameters. Moreover, the values computed by employing $\mathcal{D}_{x}$ (i.e., values in column 3) are exceptionally high, making all the software critically risky. Moreover, these values do not directly correlate to their counterparts, which demonstrates the eminence of the excluded parameters and reflects the importance of the security context of the software, reputation of developer and publisher, and penalty to assess trust-based risk level. The parameters impacting the values presented in column 3 include code dependencies, code coverage, user ratings, forks, vulnerabilities resolved, and total vulnerabilities.
@@
Table VII delineates risk scores calculated by using OWASP risk calculator by 5 assessors working in the computer systems and security domain and have been associated with research and engineering for 5 years or more. As mentioned before, results evaluated by OWASP risk calculator are highly subjective (biased) and it is evident from the scores computed by different assessors in the same setting and context. Please note that $LF$ and $IF$ represent the likelihood factor and impact factor, respectively.
@@ IX Conclusion & Future Work
In the future, the authors intend to carry out a qualitative study to rigorously assess quantitative risk assessment frameworks in the software supply chain. Another promising direction for ongoing investigation is developing a simulator for software supply chain simulation and data generation for risk assessment. Such a simulator will significantly enhance our ability to model and analyze diverse risk scenarios, enabling more comprehensive risk assessments. Through simulated data generation, we will be able to systematically test the resilience of software systems under various conditions, helping us refine risk quantification methodologies and fortify our defenses against potential vulnerabilities. Moreover, we intend to utilize techniques like reinforcement learning to facilitate the risk assessment as well as the validation process.

### v2

@@ VIII-C Evaluation & Comparisons
Table VII: Comparison - Proposed SAFER Framework Risk with/without Self-Generated Parameters.
@@
Table VIII: Risk Scores by OWASP Risk Calculator.
@@
Table VI presents 10 random samples extracted from the results generated by utilizing the OpenSSF scorecard alongside the corresponding risk scores derived from the proposed risk assessment framework SAFER for the same repositories. While comparing the outcomes yielded by both methods, it was observed that the scores computed by the two are not directly correlated. Some scores computed by the OpenSSF scorecard suggest a higher risk band as compared to the risk scores computed by SAFER, whereas some recommend a lower risk band as opposed to the calculations made by using SAFER. This indicates that the parameters or risk factors not considered by the OpenSSF scorecard have an impact on the final risk scores. This implies that the attributes taken into account by the OpenSSF scorecard to assess the security state of the software are insufficient for accurately determining the risk level associated with the software. It necessitates additional parameters to be taken into account to properly identify the risk status and make it more stringent. Please note that the bands utilized by OpenSSF scorecard for final score are low ($5\leq score\leq 10$), medium ($52.5\leq score<$ 5), high ($2.51.5\leq score<$ 2.5), critical ($score<1.5$) represented by colored rings of yellow, orange, red, and maroon respectively on the scores, whereas the risk bands applied on the final risk score in the proposed methodology are low ($<0.25$), moderate ($0.50.25\leq risk<$ 0.5), high ($0.750.5\leq risk<$ 0.75), and critical ($0.75\leq risk\leq 1$). Sample 3 has been scored as a medium risk by the OpenSSF scorecard (column 3) despite having no vulnerabilities, no dangerous workflow patterns, no defined security policy, etc. Similarly, Sample 9 has been scored as low risk by the OpenSSF scorecard even though vulnerabilities exist, the code review has less than 50% changes approved, no fuzzing or static code analysis has been used, and there is no best practices badge associated. Proper formulation instead of pre-defined values for assigning scores and severity weight to individual parameters is equally as important as selecting the most relevant influencing parameters, the developer’s and publisher’s reputation and experience, along with user experience/satisfaction. Please note that the manual verification here includes all the parameters considered by the proposed risk framework, along with the the parameters employed by the OpenSSF scorecard. Sample 1 has the following: (1) no code coverage, (2) the developer has a reputation for developing software with a significantly high number of vulnerabilities, (3) the developer has little experience in developing software using this language, (4) no fuzzing, (5) no security policy, (6) no best practices, and (7) no information on dangerous workflows, (8) high number of detected vulnerabilities, (9) no branch protection, etc. Therefore, manual verification suggests that the risk level is critical.
Similarly, Sample 8 has the following: (1) no code coverage, (2) the developer has a reputation for developing software with a moderate number of vulnerabilities, (3) the developer has little experience in developing software using this language, (4) the context of the software is other than security or automation, (5) no fuzzing, (6) no security policy, (7) no best practices, and (8) no information on dangerous workflows, etc. Accordingly, manual verification suggests that the risk level is high. Please note that the manual verification risk rating is not according to the number of quoted risk factors. It is according to the severity of the risk factors and the scoring against them.
@@
Table VII presents the comparison of the risk scores calculated using the proposed SAFER framework by first employing the entire dataset $\mathcal{D}$ and then $\mathcal{D}_{x}$ is used which excludes the parameters (i.e., regarding the date of the software, experience in language, developer’s and publisher’s experience, number of downloads, update frequency, context, and penalty/reward) for which self-generated data points were utilized. While computing the risk scores for $\mathcal{D}_{x}$, equation 11 was modified to remove the parameter for a number of downloads and normalize the value of user ratings to bring it into the range [0,1]. The values computed with all the parameters differ significantly from the ones calculated by employing only a subset of all the parameters. Moreover, the values computed by employing $\mathcal{D}_{x}$ (i.e., values in column 3) are exceptionally high, making all the software critically risky. Moreover, these values do not directly correlate to their counterparts, which demonstrates the eminence of the excluded parameters and reflects the importance of the security context of the software, the reputation of the developer and publisher, and the penalty to assess the trust-based risk level. The parameters impacting the values presented in column 3 include code dependencies, code coverage, user ratings, forks, vulnerabilities resolved, and total vulnerabilities.
@@
Table VIII delineates risk scores calculated by using the OWASP risk calculator by 5 assessors working in the computer systems and security domain and have been associated with research and engineering for 5 years or more. As mentioned before, results evaluated by the OWASP risk calculator are highly subjective (biased), and it is evident from the scores computed by different assessors in the same setting and context. Please note that $LF$ and $IF$ represent the likelihood factor and impact factor, respectively.
@@ IX Conclusion & Future Work
In the future, the authors intend to carry out a qualitative study to rigorously assess quantitative risk assessment frameworks in the software supply chain. Another promising direction for ongoing investigation is developing a simulator for software supply chain simulation and data generation for risk assessment. Such a simulator will significantly enhance our ability to model and analyze diverse risk scenarios, enabling more comprehensive risk assessments. Through simulated data generation, we will be able to systematically test the resilience of software systems under various conditions, helping us refine risk quantification methodologies and fortify our defenses against potential vulnerabilities. Moreover, we intend to utilize techniques like reinforcement learning (RL) to evaluate whether RL can provide a more effective numeric score in addition to classification as well as facilitate the validation process.
@@ X Acknowledgments
Appendix
Dataset: The parameters included in the subset $\mathcal{D}-\mathcal{D}_{x}$ are the developer, publisher, year, language, update frequency, code length, and context. For the developer, 26 unique values represented by the 26 alphabets in the English language have been randomly generated. The same is the case with publishers, and there is no connection between the developers and the publishers. For the parameter year, the values from 2016 to 2023 were randomly selected. 3 different languages have been used here, including C, Python, and Java. Update Frequency is a random number between 0 and 1. Code length is a random number between 40 and 700. These numbers are selected by taking the number of code lines from two randomly selected repositories. The context is a random number selected from the 3 possible values of 0.2, 0.3, and 0.5. These values are selected to keep the sum of all equal to 1.
Table IX: Sample Data for Example.
Code Length Developer Publisher Year Language Update Frequency Forks Downloads Unresolved Vulnerabilities Known Vulnerabilities Dependencies Rating Code Coverage Context
$l$ $j$ $x$ $t_{k}$ Language $F^{UP}$ $f$ $N^{DS}$ $V^{U}$ $V^{T}$ $D$ $G^{US}$ $C^{C}$ $C^{TXT}$
304 W Z 2018 Java 0.08424 2003 20455 135 7556 28 7153 0.99 0.2
Table IX: Sample Data for Example (Continuation).
$V^{T}$ in all the software developed by $j$ Number of all software developed by $j$ Experience of $j$ in terms of the number of years in the same language Total number of years $j$ has been developing software Number of previously developed software by $j$ in the same language Experience of $x$ in terms of the number of software published Experience of $x$ in terms of years Proportion of Unresolved Vulnerabilities Resolved Vulnerabilities
$V^{D}$ $|S^{D}|$ $E^{DS}$ $E^{DA}$ $|S^{L}|$ $E^{PX}$ $E^{PY}$ $V^{UP}$ $V^{R}$
96912 129 2 2 33 123 2 0.017867 7421
Repositories: The repositories used in Tables VI and VIII include (i) senchalabs/connect, (ii) pouchdb/pouchdb, (iii) webcomponents/webcomponentsjs, (iv) tj/commander.js, (v) thoughtbot/neat, (vi) openpgpjs/openpgpjs, (vii) webpack-contrib/sass-loader, (viii) webdriverio/webdriverio, (ix) jneen/parsimmon, and (x) rackt/redux-simple-router.
Example: The context value of 0.2 represents a security software. The values such as $V^{D}$ have been calculated considering all 9000+ samples.
$V^{UP}=\frac{V^{U}}{V^{T}}$
$V^{R}=V^{T}-V^{U}$
Equation 1:
$R^{\textup{CD}}_{S_{i},t_{k}}=28$
Equation 3:
$w^{\textup{LC}}_{t_{k}}=\frac{96912}{129}=751.2558$
Equation 2:
$R^{\textup{CS}}_{S_{i},t_{k}}=751.2558*304=228381.8$
Equation 4:
$R^{\textup{PL}}_{S_{i},t_{k}}=1-\frac{2}{2}=0$
Equation 6:
$w^{\textup{CD}}_{S_{i},t_{k}}=1-0.99=0.01$
Equation 8:
$E^{\textup{DB}}_{S_{i},t_{k}}=\frac{33}{129}=0.255814$
Equation 7:
$w^{\textup{CS}}_{S_{i},t_{k}}=e^{-0.255814}=0.77428$
Equation 9:
$w^{\textup{PL}}_{S_{i},t_{k}}=|1-(0.01+0.77428)|=0.2157$
Equation 5:
$R^{\textup{DEV}}_{S_{i},t_{k}}$ $=0.01*28+0.77428*228381.8+0.2157*0$
$=176,831.46$
Equation 10:
$R^{\textup{PB}}_{S_{i},t_{k}}=\frac{2}{123}*\frac{1}{0.08424}=0.1930$
Equation 11:
$R^{\textup{UR}}_{S_{i},t_{k}}=1-\frac{7153}{20455}=0.6503$
Equation 13:
$C^{\textup{TXT}}=0.2$
Equation 12:
$P_{S_{i},t_{k}}=1-({0.2})^{0.017867}=0.0283$
Equation 17:
$w^{\textup{DEV}}_{S_{i},t_{k}}=\frac{1}{2003}=4.99*10^{-4}$
Equation 18:
$w^{\textup{PB}}_{S_{i},t_{k}}=\frac{135+1}{7556+2}=0.0179$
Equation 16:
$w^{\textup{UR}}_{S_{i},t_{k}}=1-(4.99*10^{-4}+0.0179)=0.9815$
Equation 15:
$R^{\textup{F}}_{S_{i},t_{k}}$ $=\frac{1}{1+e^{4-0.04(0.00049\cdot 176,831.46+0.0179\cdot 0.1930+0.9815\cdot 0.6503)}}$
$=0.3755$
Equation 19:
$R^{\textup{FP}}_{S_{i},t_{k}}=R^{\textup{F}}_{S_{i},t_{k}}=0.3755$
Tests - Errors and Missing Information: Following tests related to errors and missing information against parameters have been performed, and assumptions are embedded in the functionality of the framework for added convenience.
  1. 1.
    Do not put any value for dependencies. Did you get $R^{\textup{FP}}$ without error?
  2. 2.
    Put value for code coverage greater than 1. Did you get an error?
  3. 3.
    Put value for context other than 0.2, 0.3, or 0.5. Did you get an error?
  4. 4.
    Leave any 5 values/cells empty. Did you get $R^{\textup{FP}}$ without error?
  5. 5.
    Put data for 3 developers in the form 3,5,10. Did you get $R^{\textup{FP}}$ without error?
  6. 6.
    Put a decimal value for the number of downloads. Did you get an error message?